Genent​ Industries, Inc.​ (GII), developed standard costs for direct material and direct labor. In​ 2017, GII estimated the following standard costs for one of their major​ products, the
30−gallon heavy−duty plastic container.

Budgeted quantity

Budgeted
Price

Direct materials

0.5

pounds

$20

per pound

Direct labor

0.8

hours

$11

per hour

During​ July, GII produced and sold 3,000 containers using 1,700 pounds of direct materials at an average cost per pound of $18 and 2,450 direct manufacturing labor hours at an average wage of $11.40 per hour.
The direct manufacturing labor price variance during July is​ ____.

A. $ 600.00 unfavorable
B. $ 980.00 unfavorable
C. $ 960.00 favorable
D. $ 550.00 unfavorable
Answer #1

Labor price Variance

= (SR - AR) * ALH

= (11 - 11.40) * 2450

= - 0.40 * 2450

= 980 Unfavorable